A flat surface of area 1.25 m2 is rotated through a uniform horizontal electric field of 5 N/C. What is the electric flux if the surface is (a) parallel and (b) perpendicular to the electric field?




1
Expert's answer
2022-02-23T12:11:16-0500

Solution

Area


A=1.25m^21.25m2

Electric field E=5N/C


Electric flux is given by

"\phi=EAcos\theta"

a) when parallel then angle is 0°


So cos °=1


So flux

"\phi=5*1.25*1=6.25Wb"

b) when perpendicular then angle will be 90° so cos 90°=0


Therefore flux will be zero.